OpenStudy (anonymous):

Write the standard form of the equation of each line passing through P with slope m. P= (-5,2); m =3/1

OpenStudy (anonymous):

well can you help me lol?

OpenStudy (anonymous):

Use y-y1 = m (x-x1) substitute values of m x1 and y1 and then write in Ax + By + C = 0

OpenStudy (anonymous):

can you please explain this step by step?

OpenStudy (anonymous):

First we use the equation for a line passing through a given point i.e y-y1 = m (x-x1) we are given m = 3/1 , (-5, 2) means x1= -5 and y1 = 2 substituting in above equation we get 3 y - 2 = ---- (x - (-5)) 1 this gives y - 2 = 3 (x + 5) y - 2 = 3x + 15 transposing terms on LHS to RHS 0 = 3x - y + 15 + 2 so finally we get 3x - y + 17 = 0

OpenStudy (anonymous):

are you sure this is the right equation we are to use here?

OpenStudy (anonymous):

i mean like formula?

OpenStudy (anonymous):

@marissalovescats can you take a look at this???

OpenStudy (marissalovescats):

You should get y=3x+17 for this one

OpenStudy (anonymous):

ok so how did you convert his answer to this?

OpenStudy (marissalovescats):

y-y1=m(x-x1)

OpenStudy (marissalovescats):